# A Comparative Study on Reasoning Patterns of OpenAI’s o1 Model

### v1

@@ 4.2 Analysis of the reasoning pattern of o1
  • Self-Refinement (SF). o1 assesses its reasoning process during inference to determine if there are any issues and correct any errors.
@@ 4.4 The Number of Reasoning Tokens Across Different Tasks for o1
To investigate whether the number of reasoning tokens is related to the o1’s ability, we compute the o1’s reasoning token numbers on different tasks. Besides, we also calculate the average number of tokens on the correct and incorrect samples, respectively. As shown in Fig. 3, we observe that the number of reasoning tokens for correct and incorrect samples is similar for the same task. Instead, there is a significant difference in reasoning tokens across different tasks. Specifically, for commonsense reasoning tasks (i.e., HotpotQA and Collie), the o1’s reasoning token length is relatively short. However, for more difficult tasks like Code (i.e., USACO) and Math (i.e., AIME), the model often requires a longer reasoning process to obtain the correct answer.
Figure 3: The statistics of the number of o1’s reasoning tokens on different tasks.
@@ 6 Conclusion
In this work, we explore the capabilities of OpenAI’s o1 model in tasks involving mathematics, coding, and commonsense reasoning, and compare o1 with previous Test-time Compute methods (BoN, Step-wise BoN, Self-Refine, and Agent Workflow), where the findings are as follows. First, we find that o1 achieves better results than other Test-time Compute methods across most tasks. Second, the Agent Workflow method shows significant improvements in all tasks, but the BoN, Step-wise BoN, and Self-Refine methods yield limited improvements due to their reliance on the model’s long-context instruction-following ability and the performance of the reward model. Third, we also summarize six reasoning patterns (i.e., SA, MR, DC, SF, CI, EC) of o1, and demonstrate that SR and DC are the main reasoning patterns of o1, which indicates that these reasoning patterns are crucial for improving reasoning performance. Finally, we hope our study on the reasoning patterns of o1 can guide developers and researchers in understanding the principle of o1 and facilitate the growth of foundation models.

### v2

@@ 4.2 Analysis of the reasoning pattern of o1
  • Self-Refinement (SR). o1 assesses its reasoning process during inference to determine if there are any issues and correct any errors.
@@ 4.4 The Number of Reasoning Tokens Across Different Tasks for o1
To investigate whether the number of reasoning tokens is related to o1’s ability, we developed a rule to extract o1’s reasoning tokens and computed their count across different tasks. Additionally, we calculated the average number of tokens for both correct and incorrect samples. Furthermore, to explore the relation between input prompt length and reasoning tokens length, we also calculate the average input length. As shown in Fig. 3, we observe that the number of reasoning tokens for correct and incorrect samples is similar for the same task, and there is no clear correlation between the input prompt length and the length of the reasoning tokens. Instead, there is a significant difference in reasoning tokens across different tasks. Specifically, for commonsense reasoning tasks (i.e., HotpotQA and Collie), the o1’s reasoning token length is relatively short. However, for more difficult tasks like Code (i.e., USACO) and Math (i.e., AIME), the model often requires a longer reasoning process to obtain the correct answer.
Figure 3: The statistics of the number of o1’s reasoning tokens on different tasks. ‘ALL’ represents the average length of reasoning tokens for all samples, while ‘True’ and ‘False’ show the averages for correctly and incorrectly answered samples, respectively. ‘Input’ refers to the average length of the input prompt.
@@ 6 Conclusion
In this work, we explore the capabilities of OpenAI’s o1 model in tasks involving mathematics, coding, and commonsense reasoning, and compare o1 with previous Test-time Compute methods (BoN, Step-wise BoN, Self-Refine, and Agent Workflow), where the findings are as follows. First, we find that o1 achieves better results than other Test-time Compute methods across most tasks. Second, the Agent Workflow method shows significant improvements in all tasks, but the BoN, Step-wise BoN, and Self-Refine methods yield limited improvements due to their reliance on the model’s long-context instruction-following ability and the performance of the reward model. Third, we also summarize six reasoning patterns (i.e., SA, MR, DC, SR, CI, EC) of o1, and demonstrate that SR and DC are the main reasoning patterns of o1, which indicates that these reasoning patterns are crucial for improving reasoning performance. Finally, we hope our study on the reasoning patterns of o1 can guide developers and researchers in understanding the principle of o1 and facilitate the growth of foundation models.